please help me
27. Solve
-15<=3x-4<= -3
The solution is {x|____<=x<=__________
(type and integar or a fraction)
28. Find the slope and the y-intercept.
f(x)= -4x-5
The slope is __________
The y-intercept is (0,___)
29. Solve for the indicated letter
a=7b,for b
The solution is b= ______
31. Use < or > to make the statement true.
-12 ___15
Thank you bunches in advance. You are saving my life.


Answer by JBarnum(2146)

-15<=3x-4<= -3

you seperate the 2 inequations
-15<=3x-4
3x-4<= -3
solve them like a real equality
(-11/3)<=x
x<=(1/3)
Plug them into correct spot on answer
x|(-11/3)<=x<=(1/3)
-----------------------------------------------
Find the slope and the y-intercept.
f(x)= -4x-5
They give u 0=x since y=f(x) then y=-4(0)-5= -5
to get the intercept set y to 0 solve for x:
0=-4x-5
5=-4x
-5/4=x
The slope is -5/4
The y-intercept is (0,-5)
-----------------------------------
Solve for the indicated letter
a=7b,for b ((Divide by 7))
The solution is b= a/7
------------------------------------------
Use < or > to make the statement true.
A negative number is always less than a possitive number
think of money is $15 in your pocket more then $12 u owe your mom
also look on a x-axis
-12 -11 -10 -9 -8 -7 -6 -5 -4 -3 -2 -1 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
-12 ___15